Danno is trying to decide which of two bonds to buy. Bond H is a 10 percent coupon, 10-year maturity, $1,000 par, January 1, 2000 issue paying annual interest. Bond F is a 10 percent coupon, 10-year maturity, $1,000 par, January 1, 2000 issue paying semiannual interest. The market required return for each bond is 10 percent. When using present value to determine the prices of the bonds, Danno will find that ________.
Answer #1

We need to compute the prices of each bond

Bond H

Using financial calculator

Input: FV= 1000; PMT =10%*1000=100; N = 10;I/Y=10

Solve for PV as -1000

Hence price of Bond H = $1000

BOND F

Using financial calculator

Input: FV= 1000; PMT =10%*1000/2=50; N = 10*2=20; I/Y = 5

Solve for PV as -1000

Hence price of Bond F = $1000

Hence Danno will find that both bonds have the same price and they are priced at par.